You find a stock with a beta of 1.3 that just paid a dividend of $1.45 that is expected to grow at 5%. If the risk-free rate is 3% and the market risk premium is 8%, what should be the price of the stock in four years?

A. $20.98

B. $22.03

C. $18.13

D. $41.12
Answer #1

Required rate=risk free rate+Beta*market risk premium

=3+(1.3*8)=13.4%

Current price=D1/(Required return-Growth rate)

=(1.45*1.05)/(0.134-0.05)

=$18.125

P4=Current price*(1+Growth rate)^4

=18.125*(1.05)^4

=$22.03(Approx).